I have been stuck on this math problem for days! I have had my step mom try and help as well and we have gotten nowhere..
An express train and a local train leave Townsville at 8a.m. and head for Cityville 90 miles away. The express train travels three times as fast as the local train, and arrives two hours ahead. Find the speed of each train.
The problem we run into, is there isn't enough information to answer the question, unless there is something were completely missing.
This is what I've tried:
d r t
ex 90 3x-2 90/3x-2
lo 90 x 90/x
90/3x-2 = 90/x
90x = 90(3x-2)
90x = 270x -180
-180x = -180
x=1
So I don't think the train was going one mile per hour.. But that's all I can figure out.. And this is nowhere in my algebra book.
Thank you for your time,
Lauren

Answer by checkley77(12844)

THE 1 YOU FOUND WAS THE TIME NOT THE RATE.
D=RT
FOR THE EXPRESS TRAIN:
90=RT
THE LOCAL TRAIN:
90=R/3(T+2)
SET THESE 2 EQUATION EQUAL:
RT=R/3(T+2)
RT=(RT+2R)/3 CROSS MULTIPLY
3RT=RT+2R
3RT-RT=2R
2RT=2R CANCEL THE RS.
2T=T
T=1 HOUR.
90=R*1
90=R FOR THE EXPRESS TRAIN.
PROOF:
90/3=30 MPH FOR THE LOCAL
90=30(1+2)
90=30*3
90=90